1200=40((1-(1/(1+x)^60)/x))+((1000/(1+x)^60))


find x.


1
Expert's answer
2022-02-08T16:37:07-0500

"1200=40(1- \frac 1 {(1+x)^{60} })/ x+ \frac{1000} {(1+x)^{60}}"

"1200=40- \frac {40} {x(1+x)^{60} }+ \frac{1000} {(1+x)^{60}}"

"1200-40 =\frac{1000} {(1+x)^{60}} -\frac {40} {x(1+x)^{60} }"

"1160 =\frac{1000} {(1+x)^{60}} -\frac {40} {x(1+x)^{60} }"

"1160 =\frac{1000} {(1+x)^{60}} -\frac {40} {x(1+x)^{60} }"

"x(1+x)^{60}=\frac{1000 x} {1160} -\frac {40} {1160 }"

"x(1+x)^{60}=\frac{25 x} {29} -\frac {1} {29 }"

"29x(1+x)^{60}=25x-1" ..............(1)


By using Binomial theorem, we can write "(1+x)^{60}"

"(1+x)^{60}=1 +{60}x+\frac{60(60-1)}{1!} x^2+\frac{60(60-1)(60-2)}{2!} x^3+............"

Since we need a particular answer, as progress with above series the higher power if "x" won't be effect on the value

So neglect the high power of "x", Now we can write the above binomial expression

"(1+x)^{60}=1 +{60}x"

Now we can write the equation (1) is

"29x(1+60x)=25x-1"

"29x+1740x^2 = 25x-1"

"1740x^2 +29x-25x+1=0"

"1740x^2+4x+1=0"




by using above quadratic formula


"x=\frac{-4±\sqrt{4^2-4*1740*1}}{2*1740}"


"x=\frac{-4±\sqrt{16-6960}}{3480}"


"x=\frac{-4±\sqrt{-6944}}{3480}
= \frac{-4±{83.3306i}}{3480}"


"\boxed {x=-0.001147±.02i}" Answer


if we want only positive value, Neglect the negative part...

"\boxed {x=-0.001147+02i}" Answer

or

if we want only real value then neglect the imaginary part

"\boxed {x=-.001147}" Answer